Grade 12Mechanics

If equal volumes of water and a liquid of relative density 2.6 are mixed ,the relative density of the mixture is

Shirin
7 Years agoGrade 12

1 Answer

Arun
7 Years ago
volume of water is =m/1 volume of liquid is = m/2after mixing   mass of the mixture = (2m)after mixing volue of mixture =(m+0.385m)=1.385mso density =2m/1.385m    = 1.444